# 22-large sets are sets of Bohr recurrence

### v1

@@ 1 Introduction
We say that $S\subset\mathbb{N}$ is “large” if whenever $\mathbb{N}$ is two-colored, there are arbitrarily long arithmetic progressions with common differences in $S$.
@@ 2 Coloring and Proof Outline
Before describing the proof of Theorem 1.6, we give a remark about the technique from [6] and [9] and why it will not generalize, and we explain what we need to do instead. The proof of [6] and [9] proceeds by considering a map $f$ from $\mathbb{Z}$ to $\mathbb{R}/\mathbb{Z}$ taking $n\to\{n\alpha\}$, and coloring $\mathbb{R}/\mathbb{Z}$ in two colors. In particular, if one makes the interval $[0,1/2)$ is red and the interval $[1/2,1)$ is blue then the length of a monochromatic arithmetic progression whose common difference is bounded away from 0 in $\mathbb{R}/\mathbb{Z}$ will be bounded. This does not generalize.
@@
Let $d$ be a large positive integer. Then a two-coloring of $(\mathbb{R}/\mathbb{Z})^{d}$ will contain an infinitely long arithmetic progressions with nonzero common difference $s$, so that all coordinates of $s\in(\mathbb{R}/\mathbb{Z})^{d}$ are 0 or $1/3$.
@@
Look only at the points $\{0,1/3,2/3\}^{d}$. By Hales-Jewett we can find a combinatorial line, which correspondents to a three-term and thus an infinitely long arithmetic progression. ∎
@@ 3 The Proof
Every line will meet $\Omega_{d,1/\delta}(k)$ such cells. One can see this by for instance projecting onto the $i$th coordinate. In fact, we can find $k^{O_{d,1/\delta}(1)}$ choices of $\Omega_{d,1/\delta}(k)$ cells (which we call an orbit) so that every line meets one of them, for instance by controlling the angle of the line up to a $1/k^{O_{d,1/\delta}(1)}$ error and the intercept up to a $N_{1}/k^{O_{d,1/\delta}(1)}$ error.
@@
Hence, there are only $(N_{1}/N_{0})^{O_{d,1/delta}(1)}=k^{O_{d,1/\delta}(1)}$ possible orbits, and in a random coloring each one has a chance exponentially small in $k$ of being monochromatic. So if $k$ is chosen large enough, then we will not have length $N_{2}$ arithmetic progressions that are monochromatic.
@@ 4 Further Directions
In general, a set is not “large” in the sense of [3] if it is not a set of multiple recurrence for some dynamical system. Toral rotations are the first example of a dynamical system. One could ask also about increasingly more general classes of dynamical systems such as nilsystems, skew products, and distal systems. Since every dynamical system can be built from distal systems and weak mixing systems by the Furstenberg-Zimmer structure theorem ([7] [13]), this is perhaps a potential approach to settling the $2$-large versus large conjecture in the positive.
We know how to prove this for some nilrotations. If $\alpha$ is irrational, then the set of $n$ for which the fractional part $\{n^{2}\alpha\}$ is bounded away from 0 and $1$, so the set of $n$ so that $\{n^{2}\alpha\}||_{\mathbb{R}/\mathbb{Z}}>\delta$ for some $\delta>0$ can be shown to not be $2$-large by a similar argument, using standard facts about the equidistribution of $n^{2}\alpha$ (see [5], [12]). We strongly suspect that the same holds for all nilsystems. The construction of [1] can be rephrased as coming from a skew product on $(\mathbb{R}/\mathbb{Z})^{5}$. We generally suspect that examples like the one in [1] are not $2$-large as well. However, we do not have strong opinions on what happens for all dynamical systems, i.e. on whether the $2$-large versus large conjecture of [3] is true.

### v2

@@ 1 Introduction
We say that $S\subset\mathbb{N}$ is “2-large” if whenever $\mathbb{N}$ is two-colored, there are arbitrarily long arithmetic progressions with common differences in $S$.
@@ 2 Coloring and Proof Outline
Before describing the proof of Theorem 1.6, we give a remark about the technique from [6] and [9] and why it will not generalize, and we explain what we need to do instead. The proof of [6] and [9] proceeds by considering a map $f$ from $\mathbb{Z}$ to $\mathbb{R}/\mathbb{Z}$ taking $n\to\{n\alpha\}$, and coloring $\mathbb{R}/\mathbb{Z}$ in two colors. In particular, if the interval $[0,1/2)$ is red and the interval $[1/2,1)$ blue then the length of a monochromatic arithmetic progression whose common difference is bounded away from 0 in $\mathbb{R}/\mathbb{Z}$ will be bounded. This does not generalize.
@@
Let $d$ be a sufficiently large positive integer. Then a two-coloring of $(\mathbb{R}/\mathbb{Z})^{d}$ will contain an infinitely long arithmetic progressions with nonzero common difference $s$, so that all coordinates of $s\in(\mathbb{R}/\mathbb{Z})^{d}$ are 0 or $1/3$.
@@
Consider only the points $\{0,1/3,2/3\}^{d}$. By Hales-Jewett we can find a combinatorial line, which corresponds to a three-term and thus an infinitely long arithmetic progression. ∎
@@ 3 The Proof
Every line will meet $\Omega_{d,1/\delta}(k)$ such cells in a segment that is an $\Omega_{d}(1)$ proportion of the cell’s side length. One can see this by for instance projecting onto the $i$th coordinate. In fact, we can find $k^{O_{d,1/\delta}(1)}$ choices of $\Omega_{d,1/\delta}(k)$ cells (which we call an orbit) so that every line meets one of them in an $\Omega_{d}(1)$ proportion of the cell’s side length, for instance by controlling the angle of the line up to a $1/k^{O_{d,1/\delta}(1)}$ error and the intercept up to a $N_{1}/k^{O_{d,1/\delta}(1)}$ error. This means that the arithmetic progression will meet all the cells in the orbit.
@@
Hence, there are only $(N_{1}/N_{0})^{O_{d,1/\delta}(1)}=k^{O_{d,1/\delta}(1)}$ possible orbits, and in a random coloring each one has a chance exponentially small in $k$ of being monochromatic. So if $k$ is chosen large enough, then we will not have length $N_{2}$ arithmetic progressions that are monochromatic.
@@ 4 Further Directions
In general, a set is not “large” in the sense of [3] if it is not a set of (topological) multiple recurrence for some dynamical system. Here we have handled torus rotations, which are a special kind of dynamical system. One could ask also about increasingly more general classes of dynamical systems such as nilsystems, skew products, and general distal systems. Since every dynamical system can be built from distal systems and weak mixing systems by the Furstenberg-Zimmer structure theorem ([7] [13]), this is perhaps a potential approach to settling the $2$-large versus large conjecture in the positive.
We know how to settle the $2$-large versus large conjecture for some nilrotations. If $\alpha$ is irrational, then the set of $n$ for which the fractional part $\{n^{2}\alpha\}$ is bounded away from 0 and $1$, so the set of $n$ so that $\{n^{2}\alpha\}||_{\mathbb{R}/\mathbb{Z}}>\delta$ for some $\delta>0$ can be shown to not be $2$-large by a similar argument, using standard facts about the equidistribution of $n^{2}\alpha$ (see [5], [12]). We strongly suspect that the same holds for all nilsystems. The construction of [1] can be rephrased as coming from a skew product on $(\mathbb{R}/\mathbb{Z})^{5}$. We generally suspect that examples like the one in [1] are not $2$-large as well. However, we do not have strong opinions on what happens for all dynamical systems, i.e. on whether the $2$-large versus large conjecture of [3] is true.